OpenStudy (anonymous):

Solve M = 2x + 3y for y.

OpenStudy (anonymous):

3y = M - 2x y = (M - 2x) / 3

OpenStudy (anonymous):

how dis you get that? like what are the steps?

OpenStudy (anonymous):

did*

OpenStudy (anonymous):

first we need to rearrange so that 3y is on its own you do this by subtracting 2x from both sides of the '=' sign M - 2x = 2x - 2x + 3y M - 2x = 3y now divide both sides of equation by 3 3y/3 = (M - 2x) / 3 y = (M - 2x) / 3 the brackets are there to show that all the 'M-2x' is divided by 3

OpenStudy (anonymous):

wow, thank you sooo much(: